if cos^4A-sin^4A=x, find cos^6A-sin^6A in terms of x

Grade:10

2 Answers

Sudheesh Singanamalla
13 years ago

cos^4 A - sin^4 A = x

= (cos^2 A)^2 - (sin^2 A)^2 = (cos^2 A + sin^2 A)(cos^2 A - sin^2 A) = x

==> cos^2 A - sin^2 A = x

cos^6 A - sin^6 A = (cos^2 A)^3 - (sin^2 A)^3 = (cos^2 A - sin^2 A)(cos^4 A + sin^4 A + sin^2 A. cos^2 A)

= x . ( 1 + sin^2 A .cos^2 A)

= x + x.sin^2 A.cos^2 A

 

Please approve !

 

 

vikas askiitian expert
13 years ago

cos4A-sin4A = (cos2A+sin2A)(cos2A-sin2A) = x

               = (cos2A-sin2A) = x                                                 (cos2A-sin2A = cos2A)

               cos2A = x      ..........1

              sin2A = (1-x2)1/2   ............2

 

 

we have to find the value of cos6A - sin6A = P

 P = cos6A - sin6A

    =(cos2A+sin2A)(cos4A + sin4A-cosAsinA)

    =(sin4A+cos4A-cosAsinA)

    = [ 1 + cos22A -sin2A]/2

    = [ 1+ x2 - (1-x2)1/2]/2

this is the value of expression